1.    Suppose the market demand curve for a good is Qd=1000-10p and the market supply curve is given by Qs=100+20p required.

 

i)            Calculate the equilibrium price.                             (4 Marks)

ii)          Calculate the equilibrium quantity.                        (2 Marks)

iii)        Illustrate the concept of market equilibrium.          (4 Marks)


1
Expert's answer
2021-11-09T10:49:46-0500

i) At Equilibrium Quantity supplied = Quantity Demanded

"Qs=Qd"

"1000-10P=100+20P"

"1000-100=20P+10P"

"900=30P"

"P=30"

Therefore Equilibrium Price is 30

ii) Equilibrium Quantity is gotten by use of equilibrium Price

"Qs=100+20(30)"

"Qs=100+600"

"Qs=700"

"Qd=1000-10(30)"

"Qd=1000-300"

"Qd=700"

Hence Equilibrium quantity is 700 units

iii) A market is said to be in equilibrium if the quantity supplied is equal to the quantity demanded. In this situation, the market can be said to be in equilibrium since at price of 30 the quantity supplied of 700 units is equal to the quantity demanded of 700 units too. It is equilibrium price or market clearing price and the corresponding quantity is the equilibrium quantity.